OpenStudy (anonymous):

Which is a solution to the equation 6x + y = –4? A. (2, 8) B. (1, –10) C. (2, –8) D. (1, 2)

OpenStudy (anonymous):

@Shay17

OpenStudy (matt101):

To find the answer, sub in the x-coordinate of each point into your equation, and see if the y-coordinate you calculate matches the one in the point. For instance, for the first point, (2,8), sub in the 2 for x in your equation: 6x + y = –4 6(2) + y = -4 12 + y = -4 y = -16 So for this equation, when x = 2, y = -16 and not 8. Therefore (A) cannot be the right answer. Try this for the other points and see which one works!